# The universal family of punctured Riemann surfaces is Stein

### v3

@@ Abstract.
We show that for any pair of integers $g\geq 0$ and $n\geq 1$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold. We describe its basic function theoretic properties and pose several challenging questions.
@@
(1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
@@
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$. If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [28, Theorem 5.4.3]). In this paper we prove the following result.
Theorem 1.
If $n\geq 1$ then the universal family $V(g,n)$ is a Stein manifold.
@@
This is a special case of Theorems 8 and 9 given in the sequel. The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [28, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [27]). The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1), corresponding to the punctures of $V(g,n)$, are called canonical sections. The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [25] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$. See also [26]. Earle and Kra proved in [10, p. 50] that for $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
@@
Theorem 1 has interesting consequences for function theory and complex geometry. By classical results of Remmert, Bishop, and Narasiman (see [15, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$. When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [11] and Schürmann [31], or the exposition in [15, Secs. 9.3–9.4]). Since every fibre of $V(g,n)$ is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
Problem 2.
@@
Except in the trivial cases, $T(g,n)$ is not affine algebraic, and it is not reasonable to expect that $V(g,n)$ is affine algebraic.
Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$. The following is an immediate corollary to this observation, Theorem 1, and the main result of Oka theory [15, Theorem 5.4.4]. For Oka manifolds, see [15, Chap. 5] and the surveys [13, 12, 17].
Corollary 3.
If $n\geq 1$ and $(g,n)\neq(0,1)$ then $V(g,n)$ is not simply connected, and its homotopy type is that of a finite bouquet of circles. In this case, homotopically nontrivial maps $V(g,n)\to Y$ exist whenever the manifold $Y$ is not simply connected. This gives the following corollary. The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$.
Corollary 4.
If $n\geq 1$, $(g,n)\neq(0,1)$, and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre. In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [15, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial. Since $V(g,n)$ is Stein, the Oka–Grauert principle ([19] or [15, Theorem 3.2.1]) implies the following.
Proposition 5.
@@
Corollary 6.
@@
Part (a) follows by applying Proposition 5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$. To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
@@
By the Gunning–Narasimhan theorem [23], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 6 (b), the following is a natural question.
Problem 7.
@@
Let $n\geq 1$. Does there exist a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is an immersion?
@@
By [14, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points. The problem is to find a function $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$. Note that [16, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is a holomorphic immersion. Problem 7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$. However, this is not the only problem. Since the submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$, there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
In [16] the Oka principle was established for families of maps from very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [16, Theorem 1.6] are merely continuous or smooth in the parameter. In [18] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds. Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$. This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
@@
Another interesting question is whether the Riemann surfaces $M_{t}$ in the universal Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$. For background, see [29] and [4]. By [16, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter. It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature. Each single surface in the family can be made such by [6].
@@
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$, Theorem 1 is an immediate consequence of the following result with arbitrary Stein manifold as the base. See also Theorem 9 for a more general result.
Theorem 8.
@@
Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
@@
The conclusion fails if the fibres of $\pi$ have complex dimension >1>1 or the sections $s_{i}$ are not holomorphic. In such case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$. Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [24] where the reader can find references to earlier works. In our case, $Z$ is not compact unless $X$ is a point.
Proof of Theorem 8.
@@
Let $\pi:Z\to X$ be as in the theorem. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
@@
Recall the following result (see Grauert and Remmert [21, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [21, Chap. A.3.5]). Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
@@
By a theorem of Siu [32], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [21, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [20] will then imply that $Z\setminus H$ is Stein.
@@
(2) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
@@
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [15, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H=\pi^{-1}(X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$. Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$. Set
@@
By the construction, the restriction of $\rho$ to each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X)$ is strongly subharmonic, and there is an open neighbourhood $U_{0}\subset U\subset Z$ of $H$ such that $\rho=\phi$ holds on $U_{0}\setminus H$. In particular, $\rho$ is strongly plurisubharmonic on $U_{0}\setminus H$.
Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact. Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that $\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$ is a strongly plurisubharmonic exhaustion function. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by $(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)$ for $z\in Z$ and $\xi\in T_{z}Z$, $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if $(dd^{c}\rho)(z,\xi\wedge J\xi)>0$ for every tangent vector $0\neq\xi\in T_{z}Z$. (Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since the function $\rho$ constructed above is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that

### v4

@@ Abstract.
We show that the universal Teichmüller family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold for any $n>0$. We describe its basic function theoretic properties and pose several challenging questions. We show in particular that the space of fibrewise algebraic functions on the universal family is dense in the space of holomorphic functions, and there is a fibrewise algebraic map of the universal family in a Euclidean space which is an embedding over any given relatively compact domain in the Teichmüller space. We also obtain a relative Oka principle for holomorphic fibrewise algebraic maps of the universal family to any flexible algebraic manifold.
@@ 1. Introduction
(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
@@
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$. If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [39, Theorem 5.4.3]).
In this paper, we establish several function theoretic properties of the universal family $V(g,n)$. Our first main result is the following.
Theorem 1.1.
The Teichmüller family $V(g,n)$ for $n\geq 1$ is a Stein manifold.
@@
This is a special case of Theorems 2.1 and 2.2 given in the sequel. The result is obvious if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ is biholomorphic to $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [39, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ since the Kobayashi metric and the Carathéodory metric on it differ (see Marković [38]), while they agree on a convex domain (see Lempert [37]). Note also that $\widehat{V}(g,n)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,n)$.
We provide two proofs of Theorem 1.1. The first more elementary one in Sect. 2 shows that $V(g,n)$ admits a strongly plurisubharmonic exhaustion function, hence is Stein by a theorem of Grauert [26]. The second one follows from Theorem 3.1 which shows that holomorphic fibrewise algebraic functions on $V(g,n)$ separate points and provide holomorphic convexity. The proof uses Grauert’s theorem on coherence of direct images of coherent analytic sheaves and other techniques of complex geometry.
@@
Theorem 1.1 has several interesting consequences. By classical results of Remmert, Bishop, and Narasiman (see [19, Theorem 2.4.1] and the references therein), $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$. When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [16] and Schürmann [42], or the exposition in [19, Secs. 9.3–9.4]).
Since the Teichmüller space $T(g,n)$ is contractible, the submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class. The following is a corollary to this observation, Theorem 1.1, and the main result of Oka theory [19, Theorem 5.4.4]. (See also the surveys [20, 21, 23].)
Corollary 1.2.
If $n\geq 1$ and $(g,n)\neq(0,1)$ then $V(g,n)$ is not simply connected, and its homotopy type is that of a finite bouquet of circles. In this case, homotopically nontrivial maps $V(g,n)\to Y$ exist whenever the manifold $Y$ is not simply connected. This gives the following corollary. The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$. The result obviously holds for g=0,n=1g=0,\ n=1 since $V(0,1)=\mathbb{C}$.
Corollary 1.3.
If $n\geq 1$ and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre. In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [19, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
Since every fibre of $V(g,n)$ is an affine algebraic curve, it is natural to expect that $V(g,n)$ admits holomorphic functions which are algebraic on every fibre. This is indeed the case. We show in Theorem 3.4 that the algebra $\mathscr{A}(V(g,n))$ of fibrewise algebraic functions on $V(g,n)$ is dense in the algebra $\mathscr{O}(V(g,n))$ of holomorphic functions in the compact-open topology. Furthermore, functions in $\mathscr{A}(V(g,n))$ provide a proper embedding of $V(g,n)$ in a Euclidean space over any relatively compact domain in $T(g,n)$ (see Theorem 3.1), and fibrewise algebraic maps of $V(g,n)$ to any flexible algebraic manifold satisfy a relative Oka principle (see Theorem 4.1). The following problem remains open.
Problem 1.4.
@@
The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial. Since $V(g,n)$ is Stein, the Oka–Grauert principle (see Grauert [25] or [19, Theorem 3.2.1]) implies the following.
Proposition 1.5.
@@
This fails in the algebraic category, even on a single fibre. In fact, the tangent bundle of a finitely punctured compact Riemann surface is not algebraically trivial in general.
Corollary 1.6.
@@
Part (a) follows by applying Proposition 1.5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$. To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
@@
By the Gunning–Narasimhan theorem [31], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 1.6 (b), the following is a natural question.
Problem 1.7.
@@
Let $n\geq 1$. Is there a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to every fibre of $V(g,n)$ is an immersion?
@@
By [17, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points. The problem is to find $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$. Note that [22, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre is a holomorphic immersion. Problem 1.7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 1.6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$. However, this is not the only problem. Since the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$ (see Hubbard [34, 35] and Earle and Kra [15, p. 50] for a precise description of sections of $\widehat{V}(g,n)$ and $V(g,n)$), there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
@@
Another interesting question is whether the Riemann surfaces $M_{t}$ in the Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$. For background, see [40] and [4]. By [22, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter. It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature. Each single surface in the family can be made such by [7, 6].
It is natural to wonder whether the results of this paper extend to infinite dimensional Teichmüller families. The analogous problem with continuous or smooth dependence of the complex structures and the holomorphic maps on the parameter was studied in [22] for very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t$ in a topological space $T$. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. Under mild assumptions on the family $\{J_{t}\}_{t\in T}$, $J_{t}$-holomorphic functions on $(M,J_{t})$ satisfy the Runge approximation theorem with continuous or smooth dependence on $t\in T$ [22, Theorem 1.1]. For a class of parameter spaces including finite CW complexes we also have the Oka principle for continuous or smooth families of holomorphic maps from $(M,J_{t})$ to any Oka manifold [22, Theorem 1.6]. These results were extended in [24] to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds.
2. Proof of Theorem 1.1
@@
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ (1.1), Theorem 1.1 is an immediate consequence of the following result with arbitrary Stein manifold as the base.
Theorem 2.1.
@@
Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected fibres, and $s_{1},\ldots,s_{n}:X\to Z$ $(n\geq 1)$ are holomorphic sections with pairwise disjoint images. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
@@
The assumption that the sections $s_{1},\ldots,s_{n}$ have pairwise disjoint images is inessential and is imposed only for convenience of the proof; see Theorem 2.2 for a more general result. The conclusion fails if the fibres of $\pi$ have complex dimension >1>1, or if the sections $s_{i}$ are not holomorphic. In such a case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$. Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [32] where the reader can find references to earlier works. In our case, $Z$ is not compact unless $X$ is a point.
Proof of Theorem 2.1.
@@
Let $\pi:Z\to X$ be as in the theorem. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact connected Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
@@
Recall the following result (see Grauert and Remmert [27, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [27, Chap. A.3.5]). Hence, it suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a holomorphic section $s:X\to Z$ is a Stein manifold.
@@
By a theorem of Siu [44], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold [27, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [26] will then imply that $Z\setminus H$ is Stein.
@@
(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
@@
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [19, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$. Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$. Set
@@
By the construction, the restriction of $\rho$ to each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X)$ is strongly subharmonic, and there is an open neighbourhood $U_{0}\subset U\subset Z$ of $H$ such that $\rho=\phi$ holds on $U_{0}\setminus H$. In particular, $\rho$ is strongly plurisubharmonic on $U_{0}\setminus H$. Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact. Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that
$\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$
is a strongly plurisubharmonic exhaustion function, thus proving the theorem. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by
$(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)\ \ \text{for $z\in Z$ and $\xi\in T_{z}Z$},$
the function $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if
$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{for every tangent vector $0\neq\xi\in T_{z}Z$}.$
(Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since $\rho$ is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that